# On scalable and efficient training of diffusion samplers

### v2

@@ 1 Introduction
Inference in unnormalized densities is a central challenge in machine learning, underlying probabilistic deep learning [18, 24] and many scientific applications [32, 9]. Traditionally, Markov chain Monte Carlo (MCMC) methods have been used, most prominently Metropolis-adjusted Langevin algorithms (MALA) [36] and Hamiltonian Monte Carlo (HMC) [15], but they incur repeated energy-gradient evaluations per sample. Amortized inference instead trains deep generative models to map noise to samples, enabling evaluation-free generation at test time and promising orders-of-magnitude speedups once the model is trained.
Researchers have recently focused on diffusion samplers, which parameterize continuous-time diffusion processes with neural networks, an approach inspired by successes in high-dimensional settings like image and text generation. The leading methods include flow-annealed importance sampling bootstrap (FAB) [29], generative flow networks (GFlowNets) [3], denoising diffusion samplers (DDS) [41], controlled Monte Carlo diffusion (CMCD) [42], and iterative denoising energy matching (iDEM) [1]. Because samples from the target distribution are unavailable, these samplers iterate between: (1) sample from the neural diffusion model, (2) query the energy, and (3) update the model to better match the target distribution.
Despite their promise, diffusion-based samplers struggle in high dimensions. Early in training, the neural proposal is effectively random and is not aligned with the energy landscape, leading to sample-inefficient exploration. This is in contrast with the classic training-free samplers, e.g., MALA, which leverage gradient information to steer proposals toward low-energy modes from the start. Techniques to improve the sample efficiency of diffusion samplers, like replay buffers [3] and local energy-guided refinements [20], yield only marginal gains and fail to overcome the poor quality of the initial diffusion samples. Indeed, He et al. [17] recently showed that nearly all effective neural samplers rely on Langevin parametrization, i.e., incorporating energy gradients at inference, which erodes the primary efficiency benefit of amortized sampling.
@@
We propose search-guided diffusion samplers (SGDS), a simple yet powerful framework that enables scalable and unbiased training of diffusion samplers in high-dimensional problems. A training-free Markov-chain “Searcher” explores the target density augmented with an explicit exploration reward to discover underexplored modes. The diffusion “Learner” then distills these trajectories through the trajectory balance objective [27], preserving theoretical guarantees while incorporating exploration.
@@
We show that SGDS, despite its simplicity, produces substantial gains over baseline diffusion samplers across benchmarks: classical Gaussian mixtures and the Manywell task; particle simulation problems like LJ-13 and LJ-55; and a real-world molecule, Alanine Dipeptide. Our method significantly improves sample efficiency and scalability, marking a practical path towards high-dimensional diffusion-based inference.
@@ 2.1 Diffusion samplers as controlled neural SDEs
[…]
@@
where $w_{t}$ is standard $d$-dimensional Brownian motion, $u_{\theta}$ is the drift function parameterized by $\theta$ e.g., neural networks, and $g$ is the diffusion function. The goal is to choose $\theta$ such that the terminal distribution $p_{1}^{\theta}$ induced by Equation 1 matches the target, i.e., $p_{1}^{\theta}(x)\propto R(x)$).
@@
With $T$ uniform steps of size $\Delta t:=1/T$, the SDE Equation 1 is discretized via the Euler–Maruyama scheme
@@ 2.2 Continuous GFlowNet objective for diffusion samplers
1:$\mathcal{Q}_{\text{buffer}}\leftarrow\emptyset$;  fix random target net $f_{\text{rnd}}$;  initialize predictor $\hat{f}_{\phi}$ and Learner $(P_{F}(\tau;\theta),\log Z_{\theta})$
@@
4:    ~$\tilde{\mathcal{E}}(x)\;\leftarrow\;\begin{cases}\mathcal{E}(x),&r=1,\\ \mathcal{E}(x)-\alpha\,\bigl{\|}f_{\mathrm{rnd}}(x)-\hat{f}_{\phi}(x)\bigr{\|}_{2}^{2},&r>1\end{cases}$
5:    Obtain $\{x^{(i)}_{1}\}_{i=1}^{M_{\text{chain}}}$ and $\log{\hat{Z}}$ by running $M_{\text{chain}}$ parallel MCMC for $M_{\text{iter}}$ steps on ~$\tilde{\mathcal{E}}(x)$
@@
7:    // Learner: $I$ inner iterations (even iterations: on-policy, odd iterations: off-policy)
@@
17:         $\theta\leftarrow\text{Minimize}(\mathcal{L}_{\text{TB}})$triangleright diffusion sampler update
18:         $\phi\leftarrow\text{Minimize}\Bigl{(}\tfrac{1}{|\mathcal{X}|}\!\sum_{x_{1}\in\mathcal{X}}\lVert f_{\text{rnd}}(x_{1})-\hat{f}_{\phi}(x_{1})\rVert_{2}^{2}\Bigr{)}$triangleright RND predictor update
@@
20:    Re-initialize $P_{F}(\cdot\mid\theta)$ but retain $\log Z_{\theta}$triangleright Periodic partial re-initialization
@@ 3.1 Search-guided diffusion samplers (SGDS): overall framework
Algorithmic details for each component follow in subsequent sections and Algorithm 1.
@@ 3.2 Searcher
In the initial step of the algorithm, we run $M_{\text{chain}}$ parallel chains, estimating $\log{\hat{Z}}$ which is explained in Appendix A. The Searcher then stores the collected samples in a replay buffer and passes the estimated $\log{\hat{Z}}$ to the Learner model. In subsequent rounds, we incorporate exploration uncertainty from the Learner via intrinsic rewards for exploration, modifying the Searcher’s energy landscape as:
~$\tilde{\mathcal{E}}(x)=\mathcal{E}(x)-\alpha\log r_{\text{intrinsic}}(x).$ (5)
@@ 4 Related works
Using gradient-guided MCMC for improving exploration in off-policy diffusion samplers is not new. Lemos et al. [25] employed gradient-guided MCMC to populate replay buffers for GFlowNet diffusion sampler training. Sendera et al. [38] applied parallel MALA initialized from diffusion sampler states, similar to discrete local search GFlowNet methods [20]. Our approach extends the multiple-round algorithm of Lemos et al. [25], incorporating RL techniques to boost efficiency. It can be viewed as a deeper but shorter-cycle alternative to Sendera et al. [38], whose frequent diffusion-based re-initializations overly depend on sampler performance (see comparison with TB + LS at Table 1, Table 2, and Figure 4(a)).
Leveraging Learner uncertainty to guide exploration aligns with active learning and related GFlowNet approaches [35, 21]. Following generative augmented flow network (GAFN) [35], direct injection of intrinsic reward was effective, similar to our idea (see comparison with GAFN at Table 1). While Kim et al. [21] introduced additional neural samplers called adaptive teachers (AT) as Searchers to covers high loss region it is highly unstable in large scale due to Searcher’s adversarial behavior with non-stationary objective, where our method efficiently employs MCMC-based exploration without additional neural network (see comparison with AT + LP at Table 1).
@@ 5 Experiments
In this section, [^1] Source code: https://anonymous.4open.science/r/SGDS-D38C our primary goal is to demonstrate the performance and efficiency of our proposed framework through several experiments. Specifically, we aim to showcase the sample efficiency and scalability of our method, as well as validate the effectiveness of the various training strategies we introduced. We focus on presenting results on high-dimensional tasks. In all the experiments, we use four different random seeds and average the results of each run. We provide the full experimental details and additional results in Appendix A.
@@ 5.1 Main results
[…]
@@
GAFN $320.88\pm 0.36$ $573.68\pm 29.02$ $252.80\pm 30.87$ 100M * * * 100M
AT + LP $281.56\pm 2.21$ $353.64\pm 3.48$ $72.48\pm 2.97$ 370M $462.61\pm 6.67$ $739.93\pm 4.97$ $277.32\pm 2.46$ 370M
iDEM $268.99\pm 1.22$ $414.18\pm 1.06$ $145.20\pm 1.60$ 300M $494.28\pm 2.94$ $817.32\pm 3.22$ $323.04\pm 5.69$ 300M
SGDS $320.25\pm 0.13$ $\mathbf{336.51\pm 0.11}$ $\mathbf{16.26\pm 0.22}$ 20M $614.41\pm 3.44$ $\mathbf{684.76\pm 1.30}$ $\mathbf{70.35\pm 4.31}$ 20M
@@
Table 2: ELBO, $\widehat{\text{EUBO}}$, their gap, and energy calls across Lennard-Jones potential. We denote $\widehat{\mathrm{EUBO}}$ as the EUBO metrics calculated by the reference samples provided by [1], which are not exact samples from the target distribution. Bold indicates the best performance, and * indicates large absolute values of metrics.
@@
Method ELBO $\uparrow$ $\widehat{\mathrm{EUBO}}$downarrow $\widehat{\mathrm{EUBO}}$ ELBO $\downarrow$ Energy calls ELBO $\uparrow$ $\widehat{\mathrm{EUBO}}$downarrow $\widehat{\mathrm{EUBO}}$ ELBO $\downarrow$ Energy calls
PIS $\mathbf{57.73\pm 0.22}$ $59.77\pm 0.23$ $2.04\pm 0.32$ 370K $357.19\pm 3.67$ $410.15\pm 4.90$ $45.53\pm 5.58$ 45K
TB $54.73\pm 3.02$ $67.26\pm 1.63$ $12.53\pm 3.43$ 370K * * * 45K
TB+Expl+LS $52.82\pm 0.30$ $64.81\pm 0.42$ $11.99\pm 0.52$ 3M * $563.81\pm 23.26$ 1M
iDEM $27.88\pm 5.92$ $140.25\pm 4.81$ $112.37\pm 7.63$ 300M * * * 120M
SGDS $57.68\pm 0.19$ $\mathbf{59.21\pm 0.16}$ $\mathbf{1.53\pm 0.25}$ 370K $\mathbf{363.22\pm 0.87}$ $\mathbf{396.23\pm 0.33}$ $\mathbf{33.01\pm 0.93}$ 45K
@@
[…]
@@
The baselines are primarily selected based on their strong performance demonstrated in the prior work [38], as well as their methodological relevance [35, 21] or having a different framework [1]. Specifically, iDEM [1] utilizes trajectories of length $T=1,000$ for SDE integration, whereas other baselines, including PIS [45], TB [27], AT [21], and GAFN [35], employ shorter diffusion trajectories ($T=100$) with distinct optimization objectives. We further evaluate enhanced variants of these methods incorporating LP, such as PIS+LP, TB+LP, and FL-SubTB+LP, along with exploration-enhanced (+Expl) or local search (+LS) variants introduced by Sendera et al. [38].
For the LJ potentials, we omit LP-based methods due to their poor convergence despite using additional information. We also note that our comparison with iDEM is based on our reimplementation, where we modified the hyperparameters to improve sample efficiency (see the details in Appendix A).
@@ Results.
As shown in Table 1 and Table 2, our proposed framework consistently achieves superior performance across all high-dimensional tasks (Manywell-64, Manywell-128, and LJ-55). Especially, our method demonstrates the best trade-off between performance and energy efficiency, using a small number of energy calls.
In Figure 1, one can observe that our method better captures the modes in Manywell-128 when compared to the baselines. As illustrated in Figure 2(a) and Figure 2(b), even increasing the energy budget of baselines does not allow them to surpass the performance of our proposed approach. Also, as shown in Figure 3, our framework generates high quality samples with low energy. Furthermore, for the LJ-55 potential, the distribution of interatomic distances is similar to the ground truth distribution. Additionally, our method obtains competitive results with significantly fewer samples in lower-dimensional tasks such as 40GMM, Manywell-32 (see Appendix B), and LJ-13 (see Table 2).
@@ 5.2 Ablation study
In our method, we consume energy calls during both Searcher sampling and Learner training. To evaluate the efficiency of the Learner’s on/off-policy mixing training scheme, we conduct a controlled comparison where the total energy call budget (20M) is entirely allocated to AIS on Manywell-128. We run $200$ chains on the trajectories with $T=10,000$ for AIS. As a result, even though high-reward samples were collected using AIS with much longer trajectories, the MLE Learner failed to perform amortized learning as shown in Figure 2(c).
@@
We perform an ablation study to evaluate two design components of our method when proceeding to the next training round: (1) re-initializing the Learner model, and (2) retaining the pre-trained $\log{Z}$ parameter from the previous round. Specifically, to assess the benefit of re-initialization in mitigating primacy bias, we compare our method against a fine-tuning baseline where the second-round Learner continues training from the first-round model weights without re-initialization. To isolate the effect of retaining the estimated $\log{Z}$ value, we compare against a variant where the $\log{Z}$ parameter is also re-initialized at the start of the second round. As shown in Figure 2(c), our full method outperforms both ablation variants, confirming that re-initialization is beneficial for mitigating primacy bias, and that employing the $\log{Z}$ parameter across rounds leads to better training stability and performance.
@@
We assess the effectiveness of incorporating the novelty-based intrinsic reward derived by RND [10] into the Searcher sampling process in later training rounds. In our framework, starting from the second round, the Searcher sampler drives prior samples in the direction of the target distribution and exploration signal derived from a previously trained RND module, which prioritizes underexplored modes by the Learner sampler. These dynamics guide the Searcher to focus sampling efforts on modes that remain novel and close to the target distribution across rounds. As shown in Figure 2(c), at the end of round 2, our RND-augmented approach yields a smaller EUBO–ELBO gap compared to a way of repeating the same Searcher sampling without exploration. These results demonstrate that using intrinsic rewards to adaptively bias Searcher sampling toward novel modes improves overall distributional coverage across rounds.
@@
Figure 4: Qualitative results of methods in Alanine Dipeptide. (a) Ramachandran plot consisting of two backbone torsion angles $(\phi,\psi)$ and (b) 3D visualization of generated conformations, respectively.
We also consider a real-world system, Alanine Dipeptide (ALDP), consisting of 22 atoms in vacuum at a temperature of 300K. While some previous works show promising results in sampling its conformation, they rely on low-dimensional descriptors such as rotatable torsion angles [43]. Solving ALDP at all-atom resolutions remains a challenge for existing diffusion-based neural samplers.
@@
To accurately evaluate molecular energies, we employ TorchANI [16], a PyTorch implementation of ANI deep learning potentials trained on quantum-mechanical reference data. For the Searcher, we run four parallel 55ps Langevin dynamics simulations under the TorchANI potential. In the first round, simulations are performed at 600 K to efficiently sample slow degrees of freedom; in the second round, we use 300 K to capture faster motions and collect high-reward samples. The Learner and RND models use the $E(3)$-equivariant graph neural network (EGNN) architecture [37] based on atomic coordinates. We provide details in Appendix A.
@@
To establish a baseline, we generated a reference-state ensemble via a 100 ns Langevin-dynamics simulation at 300 K. We evaluated three methods: PIS, TB, and a local-search variant of TB that employs the same Langevin dynamics as the Searcher. We compared SGDS to maximum-likelihood estimation (MLE) of forward path distribution, training MLE on 2.5 times more samples, generated by the same Searcher without RND, using 10 parallel simulations to match the total number of energy evaluations. We omit the LP methods since they have large absolute values of EUBO and ELBO. We exclude comparison with Volokhova et al. [43], as they consider only rotatable torsion angles, and with Midgley et al. [29], which employs a discrete normalizing flow on internal coordinates, whereas our method utilizes a diffusion model in atomic coordinate space.
Table 3: Comparison of ELBO, EUBO, and energy calls on Alanine Dipeptide (ALDP). Bold indicates the best performance, and * indicates large absolute values of metrics.
Method (Energy calls) ELBO [×$\times 10^{3}$] $\uparrow$ EUBO [×$\times 10^{3}$] $\downarrow$
PIS (1M) $516.912\pm 5.357$ $601.565\pm 87.771$
TB (1M) * *
TB+Expl+LS (3M) $519.614\pm 0.015$ $538.479\pm 0.198$
MLE (1M) $520.618\pm 0.142$ $538.032\pm 0.000$
SGDS (1M) $\mathbf{520.916\pm 0.165}$ $\mathbf{538.025\pm 0.003}$

### v4

@@ 1 Introduction
Inference in unnormalized densities is a central challenge in machine learning, underlying probabilistic deep learning [19, 25] and many scientific applications [33, 9]. Traditionally, Markov chain Monte Carlo (MCMC) methods have been used, most prominently Metropolis-adjusted Langevin algorithms (MALA) [37] and Hamiltonian Monte Carlo (HMC) [15], but they incur repeated energy‐gradient evaluations per sample. Amortized inference instead trains deep generative models to map noise to samples, enabling evaluation‐free generation at test time and promising orders‐of‐magnitude speedups once the model is trained.
Researchers have recently focused on diffusion samplers, which parameterize continuous‐time diffusion processes with neural networks, an approach inspired by successes in high‐dimensional settings like image and text generation. The leading methods include flow‐annealed importance sampling bootstrap (FAB) [30], generative flow networks (GFlowNets) [3], denoising diffusion samplers (DDS) [43], controlled Monte Carlo diffusion (CMCD) [44], and iterative denoising energy matching (iDEM) [1]. Because samples from the target distribution are unavailable, these samplers iterate between: (1) sample from the neural diffusion model, (2) query the energy, and (3) update the model to better match the target distribution.
Despite their promise, diffusion‐based samplers struggle in high dimensions. Early in training, the neural proposal is effectively random and is not aligned with the energy landscape, leading to sample‐inefficient exploration. This is in contrast with the classic training-free samplers, e.g., MALA, which leverage gradient information to steer proposals toward low‐energy modes from the start. Techniques to improve the sample efficiency of diffusion samplers, like replay buffers [3] and local energy‐guided refinements [21], yield only marginal gains and fail to overcome the poor quality of the initial diffusion samples. Indeed, He et al. [18] recently showed that nearly all effective neural samplers rely on Langevin parametrization, i.e., incorporating energy gradients at inference, which erodes the primary efficiency benefit of amortized sampling.
@@
We propose search-guided diffusion samplers (SGDS), a simple yet powerful framework that enables scalable and unbiased training of diffusion samplers in high-dimensional problems. A training‐free Markov‐chain “Searcher” explores the target density augmented with an explicit exploration reward to discover underexplored modes. The diffusion “Learner” then distills these trajectories through the trajectory balance objective [28], preserving theoretical guarantees while incorporating exploration.
@@
We show that SGDS, despite its simplicity, produces substantial gains over baseline diffusion samplers across benchmarks: classical Gaussian mixtures and the Manywell task; particle simulation problems like LJ-13 and LJ-55; and real-world molecules, Alanine Di-, Tri-, and Tetra-peptide. Our method significantly improves sample efficiency and scalability, marking a practical path towards high-dimensional diffusion-based inference.
@@ 2.1 Diffusion samplers as controlled neural SDEs
Let $\mathcal{E}\colon\mathbb{R}^{d}\!\to\!\mathbb{R}$ be an energy function defining an unnormalized density $R(x)=\exp\bigl(-\mathcal{E}(x)\bigr)$. Sampling from the corresponding Boltzmann distribution $p_{\text{target}}(x)=R(x)/Z$, with partition function $Z=\int R(x)\,dx$, can be formulated as controlling the stochastic differential equation (SDE)
@@
where $w_{t}$ is standard $d$-dimensional Brownian motion, $u_{\theta}$ is the drift function parameterized by $\theta$ e.g., neural networks, and $g$ is the diffusion function. The goal is to choose $\theta$ such that the terminal distribution $p_{1}^{\theta}$ induced by Equation˜1 matches the target, i.e., $p_{1}^{\theta}(x)\propto R(x)$).
@@
With $T$ uniform steps of size $\Delta t:=1/T$, the SDE Equation˜1 is discretized via the Euler–Maruyama scheme
@@ 2.2 Continuous GFlowNet objective for diffusion samplers
[…]
@@
4:  $\tilde{\mathcal{E}}(x)\;\leftarrow\;\begin{cases}\mathcal{E}(x),&r=1,\\ \mathcal{E}(x)-\alpha\,\bigl\|f_{\mathrm{rnd}}(x)-\hat{f}_{\phi}(x)\bigr\|_{2}^{2},&r>1\end{cases}$
5:  Obtain $\{x^{(i)}_{1}\}_{i=1}^{M_{\text{chain}}}$ and $\log{\hat{Z}}$ by running $M_{\text{chain}}$ parallel MCMC for $M_{\text{iter}}$ steps on $\tilde{\mathcal{E}}(x)$
@@
7:  // Learner: II inner iterations (even iterations: on-policy, odd iterations: off-policy)
@@
17:    $\theta\leftarrow\text{Minimize}(\mathcal{L}_{\text{TB}})$ $\triangleright$ diffusion sampler update
18:    $\phi\leftarrow\text{Minimize}\Bigl(\tfrac{1}{|\mathcal{X}|}\!\sum_{x_{1}\in\mathcal{X}}\lVert f_{\text{rnd}}(x_{1})-\hat{f}_{\phi}(x_{1})\rVert_{2}^{2}\Bigr)$ $\triangleright$ RND predictor update
@@
20:  Re-initialize $P_{F}(\cdot\mid\theta)$ but retain $\log Z_{\theta}$ $\triangleright$ Periodic partial re-initialization
@@ 3.1 Search-guided diffusion samplers (SGDS): overall framework
Algorithmic details for each component follow in subsequent sections and Algorithm˜1.
@@ 3.2 Searcher
In the initial step of the algorithm, we run $M_{\text{chain}}$ parallel chains, estimating $\log{\hat{Z}}$ which is explained in Appendix˜A. The Searcher then stores the collected samples in a replay buffer and passes the estimated $\log{\hat{Z}}$ to the Learner model. In subsequent rounds, we incorporate exploration uncertainty from the Learner via intrinsic rewards for exploration, modifying the Searcher’s energy landscape as:
$\tilde{\mathcal{E}}(x)=\mathcal{E}(x)-\alpha\log r_{\text{intrinsic}}(x).$ (5)
@@ 4 Related works
Using gradient-guided MCMC for improving exploration in off-policy diffusion samplers is not new. Lemos et al. [26] employed gradient-guided MCMC to populate replay buffers for GFlowNet diffusion sampler training. Sendera et al. [39] applied parallel MALA initialized from diffusion sampler states, similar to discrete local search GFlowNet methods [21]. Our approach extends the multiple-round algorithm of Lemos et al. [26], incorporating RL techniques to boost efficiency. It can be viewed as a deeper but shorter-cycle alternative to Sendera et al. [39], whose frequent diffusion-based re-initializations overly depend on sampler performance (see comparison with TB + LS at Table˜1, Table˜2, and Figure˜4(a)).
Leveraging Learner uncertainty to guide exploration aligns with active learning and related GFlowNet approaches [36, 22]. Following generative augmented flow network (GAFN) [36], direct injection of intrinsic reward was effective, similar to our idea (see comparison with GAFN at Table˜1). While Kim et al. [22] introduced additional neural samplers called adaptive teachers (AT) as Searchers to covers high loss region it is highly unstable in large scale due to Searcher’s adversarial behavior with non-stationary objective, where our method efficiently employs MCMC-based exploration without additional neural network (see comparison with AT + LP at Table˜1).
@@ 5 Experiments
In this section, [^1] Source code: https://github.com/minkyu1022/SGDS our primary goal is to demonstrate the performance and efficiency of our proposed framework through several experiments. Specifically, we aim to showcase the sample efficiency and scalability of our method, as well as to validate the effectiveness of the various training strategies we introduced. We focus on presenting results on high-dimensional tasks. In all the experiments, we use four different random seeds and average the results of each run. We provide details of experimental settings in Section˜A.4, Table˜4, and Table˜5, and additional results in Appendix˜B.
@@ 5.1 Main results
Method ELBO $\uparrow$ EUBO $\downarrow$ $\text{EUBO}-\text{ELBO}$ $\downarrow$ Energy calls ELBO $\uparrow$ EUBO $\downarrow$ $\text{EUBO}-\text{ELBO}$ $\downarrow$ Energy calls
@@
GAFN [36] $320.88\pm 0.36$ $573.68\pm 29.02$ $252.80\pm 30.87$ 100M * * * 100M
AT [22] + LP $281.56\pm 2.21$ $353.64\pm 3.48$ $72.48\pm 2.97$ 370M $462.61\pm 6.67$ $739.93\pm 4.97$ $277.32\pm 2.46$ 370M
iDEM [1] $268.99\pm 1.22$ $414.18\pm 1.06$ $145.20\pm 1.60$ 300M $494.28\pm 2.94$ $817.32\pm 3.22$ $323.04\pm 5.69$ 300M
[…]
@@
Table 2: The gap between ELBO and $\widehat{\text{EUBO}}$, (equivariant) position Wasserstein-2, energy histogram Wasserstein-2, and energy calls across Lennard-Jones potential. We denote $\widehat{\mathrm{EUBO}}$ as the EUBO metrics calculated by the reference samples provided by [1], which are not exact samples from the target distribution. For iDEM, we reproduce the results with the hyperparameter setting which can be found in Section˜A.4. The results of Adjoint sampling are obtained from [17]. Bold indicates the best performance, * indicates large absolute values of metrics, - indicates inaccessible value from their papers, and "Div." indicates divergent training.
@@
Method $\widehat{\mathrm{EUBO}} -$ ELBO $\downarrow$ $\mathcal{W}_{2}\downarrow$ $E(\cdot)\mathcal{W}_{2}\downarrow$ Energy calls $\widehat{\mathrm{EUBO}} -$ ELBO $\downarrow$ $\mathcal{W}_{2}\downarrow$ $E(\cdot)\mathcal{W}_{2}\downarrow$ Energy calls
PIS $2.04\pm 0.32$ $1.57\pm 0.03$ $2.10\pm 0.46$ 370K $45.53\pm 5.58$ $4.28\pm 0.05$ $40.63\pm 13.68$ 45K
TB $12.53\pm 3.43$ $1.71\pm 0.07$ $10.99\pm 3.14$ 370K Div. Div. Div. 45K
TB+Expl+LS $11.99\pm 0.52$ $2.02\pm 0.04$ $18.69\pm 4.49$ 3M * $12.46\pm 0.19$ $123.02\pm 21.72$ 1M
iDEM [1] $112.37\pm 7.63$ $4.27\pm 0.02$ $39.92\pm 3.24$ 300M * $17.17\pm 0.32$ $210.87\pm 6.26$ 120M
Adjoint Sampling [17] - $1.67\pm 0.01$ $2.40\pm 1.25$ 1M - $4.50\pm 0.05$ $58.04\pm 20.98$ 1M
SGDS $\mathbf{1.53\pm 0.25}$ $\mathbf{1.56\pm 0.04}$ $\mathbf{0.89\pm 0.19}$ 370K $\mathbf{33.01\pm 0.93}$ $\mathbf{4.25\pm 0.08}$ $\mathbf{32.17\pm 12.90}$ 45K
@@
In this work, we compare the performance of our proposed framework against baselines on multiple benchmark tasks, including 40GMM, Manywell-32/64/128, LJ-13, and LJ-55. We evaluate methods using three metrics: the Evidence Lower Bound (ELBO), Evidence Upper Bound (EUBO) [7], and the $\text{EUBO}-\text{ELBO}$ gap. A smaller gap between ELBO and EUBO indicates a more accurate approximation of the target distribution.
@@
For the manywell potential, the baselines are primarily selected based on their strong performance demonstrated in prior work [39], as well as their methodological relevance [36, 22] or their different framework [1]. Specifically, iDEM [1] utilizes trajectories of length $T=1,000$ for SDE integration, whereas other baselines, including PIS [47], TB [28], AT [22], and GAFN [36], employ shorter diffusion trajectories ($T=100$) with distinct optimization objectives. We further evaluate enhanced variants of these methods incorporating LP, such as PIS+LP, TB+LP, and FL-SubTB+LP, along with exploration-enhanced (+Expl) or local search (+LS) variants introduced by Sendera et al. [39]. We describe the details of the abbreviations related to baselines in Section˜A.1. For the LJ potentials, we compare against Adjoint Sampling [17], as well as iDEM, PIS, TB, and its variant TB+Expl+LS. We omit LP-based methods due to their divergent training.
@@ Results.
As shown in Table˜1 and Table˜2, our proposed framework consistently achieves superior performance across all high-dimensional tasks: Manywell-64, Manywell-128, and LJ-55. Especially, our method demonstrates the best trade-off between performance and efficiency of energy call.
In Figure˜1, one can observe that our method better captures the modes in Manywell-128 when compared to the baselines. As illustrated in Figure˜2(a) and Figure˜2(b), even increasing the energy budget of baselines does not allow them to surpass the performance of our proposed approach. Also, as shown in Figure˜3, our framework generates high quality samples with low energy. Furthermore, for the LJ-55 potential, the distribution of interatomic distances is similar to the ground truth distribution. Additionally, our method obtains competitive results with significantly fewer samples in lower-dimensional tasks such as 40GMM, Manywell-32 (see Section˜B.1), and LJ-13 (see Table˜2).
@@ 5.2 Ablation study
In our method, we consume energy calls during both Searcher sampling and Learner training. To evaluate the efficiency of the Learner’s on/off-policy mixing training scheme, we conduct a controlled comparison where the total energy call budget (20M) is entirely allocated to AIS on Manywell-128. We run $200$ chains on the trajectories with $T=10,000$ for AIS. As a result, even though high-reward samples were collected using AIS with much longer trajectories, the MLE Learner failed to perform amortized learning as shown in Figure˜2(c).
@@
We perform an ablation study to evaluate two design components of our method when proceeding to the next training round: (1) re-initializing the Learner model, and (2) retaining the pre-trained $\log{Z}$ parameter from the previous round. Specifically, to assess the benefit of re-initialization in mitigating primacy bias, we compare our method against a fine-tuning baseline where the second-round Learner continues training from the first-round model weights without re-initialization. To isolate the effect of retaining the estimated $\log{Z}$ value, we compare against a variant where the $\log{Z}$ parameter is also re-initialized at the start of the second round. As shown in Figure˜2(c), our full method outperforms both ablation variants, confirming that re-initialization is beneficial for mitigating primacy bias, and that employing the $\log{Z}$ parameter across rounds leads to better training stability and performance.
@@
We assess the effectiveness of incorporating the novelty-based intrinsic reward derived by RND [10] into the Searcher sampling process in later training rounds. In our framework, starting from the second round, the Searcher sampler drives prior samples in the direction of the target distribution and exploration signal derived from a previously trained RND module, which prioritizes underexplored modes by the Learner sampler. These dynamics guide the Searcher to focus sampling efforts on modes that remain novel and close to the target distribution across rounds. As shown in Figure˜2(c), at the end of round 2, our RND-augmented approach yields a smaller EUBO–ELBO gap compared to a way of repeating the same Searcher sampling without exploration. These results demonstrate that using intrinsic rewards to adaptively bias Searcher sampling toward novel modes improves overall distributional coverage across rounds.
@@
Figure 4: Qualitative results of methods in three peptides. (a) Ramachandran plot of Alanine Dipeptide with two backbone torsion angles $(\phi,\psi)$ and (b) 3D visualization of generated conformations.
We also consider three real-world systems, Alanine Di-, Tri-, and Tetra-peptide, consisting of 23, 33, 43 atoms in vacuum at a temperature of 300K. While some previous works show promising results in sampling conformation of Alanine Dipeptide, they rely on low-dimensional descriptors such as rotatable torsion angles [45]. Solving these three peptides at all-atom resolutions remains a challenge for existing diffusion-based neural samplers.
@@
To accurately evaluate molecular energies, we employ TorchANI [16], a PyTorch implementation of ANI deep learning potentials trained on quantum‐mechanical reference data. For the Searcher, we run four parallel 55ps Langevin dynamics simulations under the TorchANI potential. In the first round, simulations are performed at 600 K to efficiently sample slow degrees of freedom; in the second round, we use 300 K to capture faster motions and collect high-reward samples. The Learner and RND models use the $E(3)$-equivariant graph neural network (EGNN) architecture [38] based on atomic coordinates. We provide details in Section˜A.4.
@@
We compare our method with maximum‐likelihood estimation (MLE) as well as diffusion-based neural samplers: PIS, TB, and TB+Expl+LS. We generate reference ensemble using 100 ns Langevin‐dynamics simulation at 300K. In MLE, the likelihood of forward path distribution is maximized over the samples from backward path distribution. To match the number of energy calls of MLE with our method, we collect 2.5 times more samples than our method, using the same searcher as our method without RND. TB+Expl+LS utilizes the same langevin dynamics for local search algorithm. We omit the LP methods since they have large absolute values of EUBO and ELBO. We exclude comparison with Volokhova et al. [45], as they consider only rotatable torsion angles, and with Midgley et al. [30], which employs a discrete normalizing flow on internal coordinates, whereas our method utilizes a diffusion model in atomic coordinates.
Table 3: EUBO-ELBO gaps on peptides in units of $10^{3}$. Bold indicates the best performance. * indicates large absolute values of metrics, and "Div." indicates divergent training.
Method Dipeptide Tripeptide Tetrapeptide
PIS (1M) $84.65\pm 9.93$ $35.23\pm 2.43$ *
TB (1M) Div. Div. Div.
TB+Expl+LS (3M) $18.86\pm 0.19$ $6.41\pm 0.34$ Div.
MLE (1M) $17.41\pm 0.14$ $6.11\pm 0.02$ $29.07\pm 0.56$
SGDS (1M) $\mathbf{17.11\pm 0.16}$ $\mathbf{5.60\pm 0.03}$ $\mathbf{26.60\pm 0.21}$